MONT 105N – Analyzing Environmental Data Paper Assignment – Views of Nature in Art due: March 1, 2019

This assignment is based on several still life paintings and one mosaic at the Worcester Art Museum that you should find in our visit on Thursday, February 24:

1. Crossbow and Dead Game, by Jan Fyt (Flemish, mid-17th century) – on view in Gallery 204

2. Still Life with Blue Vase and Fruit, by Luigi Lucioni (American, 1924-1926) – on view in Gallery 421

3. Banquet Still Life, by Abraham van Beyeren (Dutch, 1655) – on view in Gallery 204 4. “Worcester Hunt Floor Mosaic”, from Antioch, Roman, early 6th century CE) – on

view in Gallery 102

(Note: You can find electronic image files of all of these online if you want to refresh your memories about them and/or check details. But seeing a small electronically-rendered photograph is no substitute for the actual experience of seeing the work in its true scale with your own eyes. If you do view them online, please include the URL and the date you downloaded the image in your list of sources consulted.)

Assignment

You should compose an essay addressing all of the following questions.

1. Is any of these a portrayal of the natural world as it actually is? What aspect of human (or other) interaction with the natural world do we see in each of them?

2. (“Compare and contrast”) Can you identify a specific view of the relationship between humans and the natural world in each of these works? Are some different from others? Are some similar to others? In what ways? What choices (both in the subjects the artists chose to portray, and in how those subjects are portrayed, the techniques used, etc.) have the artists made to highlight those views? Discuss any connections with the Bryson Rhopography chapter that you can identify.

3. Which of these artworks made the biggest impression on you and why? Taken purely as an example of painting, I think many art critics and art historians would probably say that Crossbow and Dead Game by Jan Fyt is superior to the other paintings. Would you agree? Why or why not? On the other hand, mosaic is a completely different medium. Is it fair to make a comparison in this case?

Most of these questions are not asking solely for a personal opinion. So you should try as much as possible to justify what you are saying by citing details from or about the paintings as evidence.

Important Note: It is perfectly OK with me if you want to share a rough draft of your essay with other students in the class, or with me, to get comments and suggestions. “Testing out” your ideas that way is often an important part of getting them down on paper in an effective form.

Prepare your essay using Google Docs, Microsoft Word or similar software and submit as a file attachment to this email address:

                           jlittle@holycross.edu

Your assignment should arrive in my inbox no later than 5:00pm on Friday, March 1. Aim for around 3 double-spaced pages. (Going a bit short or long is OK, though, as long as you have addressed the questions above adequately and succinctly.)

For this assignment, your are not expected or required to refer to any sources other than the excerpt from the Bryson chapter that we discussed in class. But if you do, please include a complete listing of all sources you consulted in a numbered References section at the end. If you take a direct quotation from one of them, identify it with an in-line reference of the form “(author, page)” where the “author” refers to the author of an item in the list of references at the end. Include a page number for printed sources. Please do not use online sources for this assignment.